OpenStudy (anonymous):

a.) Find the Taylor polynomial of degree 3, expanded about a=0, for f(x)= sin(2x). I got P3(x)= 2x-8/3!(x)^3, which is correct. What I don't understand is this part: b.) Use your answer from part a.) to approximate sin(2/3) So to do this I just plugged in 2/3 into x, like this: sin(2/3)= (2/3)-8/3!(2/3)^3 BUT my instructor crossed out the 2's on top and replaced it with 1's. Why is that?

OpenStudy (anonymous):

\[P _{3}^{}(x)= 2x-8/3!(x)^3\]

OpenStudy (anonymous):

\[\sin(2/3)= 2(2/3)-8/3!(2/3)^3\]

OpenStudy (anonymous):

The Taylor polynomial you found is the expansion of sin(2x). If you want sin(2/3), you need to replace substitute x= 1/3.

OpenStudy (anonymous):

you approximated sin(2x) so x=1/3 to approximate sin(2/3)